OpenStudy (anonymous):

find the missing number 48/88=?/11

OpenStudy (anonymous):

48/8 = ?

OpenStudy (anonymous):

48/88=?/11 using cross multiplication: 48*11=?*88 528=?*88 528/88=? 6=?